Find three consecutive odd intergers such that the sum of the smaller two is three times the largest increased by seven
Answer by addingup(3677)

Let the three consecutive odd integers be x, x+2, and x+4
x+x+2 = 3(x+4)+7
2x+2 = 3x+12+7
2x+2 = 3x+19
-x = 17
x = -17 this is the smallest number
Next is -17+2 = -15
Next is -15+2 = -13
Keep in mind that the further you go into negative territory the smaller the number is, so -100 is smaller than -0.01
----------------------------
Check:
Our numbers are: -17, -15, -13
the sum of the smaller two is 3 times the largest +7
-17+-15 = 32 this is the sum of the smaller two
-13(3)+7 = 32 this is the largest of the numbers times 3 +7
as you see the answer in both cases is 32, so this answer is correct.